# ./references.bib

### v1

@@ 2 Pseudospectrum for joint approximate eigenstates
Equation (5) is taken to be the quadratic composite operator in Section 1 of [cerjan_loring_vides_2023]. After some properties are proven about the smallest eigenvalue of the quadratic composite operator, the quadratic gap function is defined.
Given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ where each of the $A_{i}$ are Hermitian operators and probe site $\bm{\lambda}$ in $\mathbb{R}^{d+1}$, the quadratic gap is defined to be
@@
The square root is a consequence of proving this quantity is related to the smallest singular value of a related matrix $M_{\bm{\lambda}}(\bm{A})$ which is the topic of Proposition II.1 in [cerjan_loring_vides_2023]. Note that by construction, the quadratic composite operator has real eigenvalues and is positive semidefinite. Thus the quadratic gap is well defined. Informaly a key idea with this gap function is that
$\mu_{\bm{\lambda}}^{\text{C}}(\bm{A})=\sigma_{\min}\left(\left(L_{\bm{\lambda}}\left(\bm{A}\right)\right)^{2}\right)\approx\sigma_{\min}\left(\text{Q}_{\bm{\lambda}}\left(\bm{A}\right)\right)\leq\mu_{\bm{\lambda}}^{\text{Q}}\left(\bm{A}\right)$
when $\mu_{\bm{\lambda}}^{Q}(\bm{A})\leq 1$. This also provides a corresponding state $\bm{\psi}$ that is an approximate eigenstate of the constituent observables/matrices. For a Comparison between the spectral localizer gap and quadratic gap function, see Proposition II.4 in [cerjan_loring_vides_2023]. The main idea is that the Clifford and quadratic gaps are close if the commutators $[A_{i},A_{k}]$ are small, in other words the matrices $A_{i}$ nearly pairwise commute.
3 Clifford composite operator
As mentioned in the introduction, a property of photonic systems is that they exhibit non-Hermitian Hamiltonians. This motivates the question: What happens when $A_{d+1}=B$ is non-Hermitian in the construction of the spectral localizer and in the quadratic composite operator? Or in a physical system setting, what happens when we have a non-Hermitian Hamiltonian $H$ in the construction of the corresponding composite operators? We take inspiration from the constructions and definitions in [cerjan_koekebier_Schulz_Baldes] to motivate this Section. In the process we also introduce a new gap function and explain some of it’s properties.
One option for a construction of an equivalent spectral localizer is to consider a matrix decomposition of $B$ into the Hermitian part $B_{\textup{H}}$ and it’s skew Hermitian part $B_{\textup{S}}$ shown in Equation (6)
$B=B_{\textup{H}}+B_{\textup{S}},\quad B_{\textup{H}}=(B+B^{\dagger})/2,\quad B_{\textup{S}}=-i(B-B^{\dagger})/2.$ (6)
Both $B_{\textup{H}}$ and $B_{\textup{S}}$ are Hermitian, and thus we can use these together with position observables in the Clifford composite operator, and thus we would be able to find a unit state $\bm{\psi}$ that is approximately localized in position and energy given the theory above. In particular we could construct
@@ 3 Clifford composite operator
However this goes against the intuition that the Clifford composite operator should have similar properties as the Hamiltonian, in this case hermiticity or lack thereof.
Also, nothing necessarily stops us from using $B$ in the construction of the spectral localizer in [loring] and [cerjan_loring_vides_2023] with complex $E$. In other words $L_{(\bm{x},E)}(\bm{X},B)$. This makes it so that the resulting operator is not only not Hermitian, It is also not normal. We will be following the work in [cerjan_koekebier_Schulz_Baldes] where a special case of the non-Hermitian spectral localizer which we will refer to as the Clifford composite operator is utilized.
In the construction of the spectral localizer in Section I of [cerjan_loring_vides_2023], an arbitrary Clifford representation can be chosen. However, in the following construction for the Clifford composite operator we use a specific Clifford representation. The construction of said Clifford representation is as follows. Note that $\sigma_{x},\sigma_{y},\sigma_{z}$ are the Pauli spin matrices. For $d=2$ we use $\sigma_{x}$ and $\sigma_{y}$ and for $d=1$ we just drop $\sigma_{y}$. For $d=3$ we use $\Gamma_{1}=\sigma_{x},\Gamma_{2}=\sigma_{y},\Gamma_{3}=\sigma_{z}$. The Construction for $d=5$ is $\Gamma_{1}=\sigma_{x}\otimes\sigma_{x}$, $\Gamma_{2}=\sigma_{y}\otimes\sigma_{x}$, $\Gamma_{3}=\sigma_{z}\otimes\sigma_{x}$, $\Gamma_{4}=I_{2}\otimes\sigma_{y}$, $\Gamma_{5}=I_{2}\otimes\sigma_{z}$. Then suppose that $\{\gamma_{i}\}_{i=1}^{d}$ represents a Clifford representation for $d$ odd, we construct a Clifford representation for $d+2$ by computing
@@
If we want the Clifford representation for $d$ even, we just construct the $d+2$ Clifford representation and discard $I_{m}\otimes\sigma_{y}$. We now proceed with the definition of the Clifford composite operator as follows.
@@
Given $\bm{A}$ in $\Pi_{i=1}^{d}\mathcal{M}_{n}(\mathbb{C})$ where each of the $A_{i}$ are Hermitian operators, a non-Hermitian matrix $B$, and a probe site $(\bm{\lambda},\nu)$ in $\mathbb{R}^{d}\oplus\mathbb{C}$ we define the Clifford composite operator to be
@@
Again, this is a special case of the spectral localizer in [cerjan_koekebier_Schulz_Baldes]. From here we can manually compute and determine that the Clifford composite operator is not normal. This can be seen below when we compute the product
@@
Despite the Clifford composite operator being non-normal, we do have a property that is just as useful if not better. If $B$ is close to being Hermitian then so too is the Clifford composite operator. Formally we have that for all $\epsilon>0$, $\left\lVert B-B^{\dagger}\right\rVert<\epsilon$ implies that if we label $L=L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)$
@@
$<\epsilon+2\textrm{Im}(\nu).$
This also tells us that the closeness of the Clifford composite operator to Hermitian is also dependent on how big of a value we probe in energy. In a way this is surprising from the point of view that if a matrix $A$ is Hermitian, then it follows that $A$ is normal, indeed $A=A^{\dagger}$ implies $A^{\dagger}A=AA=AA^{\dagger}$. Here we have shown that we can have a matrix that is non-normal but close to being Hermitian, showing the vastness of the set of complex $n$ by $n$ matrices. To define the gap function for the Clifford composite operator, we take the definition in Section 2 of [cerjan_koekebier_Schulz_Baldes] that is used while studying line gapped systems. In particular it is called the spectral localizer gap, but we will refer to it by a different name.
@@
From a topological point of view, especially for line gapped systems, it seems to be instrumental to not look at $\min|\textup{Spec}(A)|$ but instead to flatten the spectrum and instead look at $\min|\textrm{Re}(\textup{Spec}(A))|$. When we are working with non-Hermitian Systems, we have to take into consideration both point gapped and line gapped systems. For this reason we propose the following definition for point gapped systems. By numerical computations in Section 7 we will demonstrate that the smallest singular value of the Clifford composite operator detects point gaps.
@@
$\dot{\mu}_{\bm{\lambda}}^{\text{C}}\left(\bm{A},B\right)=\sigma_{\min}\left(L_{\bm{\lambda}}\left(\bm{A},B\right)\right).$
One immediate consequence of the definition of the clifford radial gap is Lipschitz continuity of the Clifford radial gap in a generalized taxi-cab metric. Lipschitz continuity implies that the Clifford radial gap is differentiable almost everywhere. This also means that when trying to generate a grid size for numerical simulations, it can be determined ahead of time based on the level of precision desired between data points.
@@
$\left|\dot{\mu}_{\bm{\lambda}}^{\text{C}}\left(\bm{A},B\right)-\dot{\mu}_{\bm{\alpha}}^{\text{C}}\left(\bm{C},D\right)\right|\leq\left(\sum_{i=1}^{d}\left\lVert A_{i}-C_{i}\right\rVert+\left\lVert\lambda_{i}-\alpha_{i}\right\rVert\right)+\left\lVert B-D\right\rVert+\left\lVert\nu-\beta\right\rVert.$
@@
$\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)-L_{(\bm{\alpha},\beta)}\left(\bm{C},D\right)\right\rVert$ $\leq\sum_{i=1}^{d}\left\lVert\left((A_{i}-C_{i})+(\lambda_{i}-\alpha_{i})I_{2m}\right)\otimes\Gamma_{i}\right\rVert$
$\hskip 21.68121pt+\left\lVert\left((B-D)+(\nu-\beta)I_{m}\right)\otimes\begin{bmatrix}I_{m}&\\ &0_{m}\end{bmatrix}\right\rVert$
$\leq\left(\sum_{i=1}^{d}\left\lVert(A_{i}-C_{i})+(\lambda_{i}-\alpha_{i})I_{2m}\right\rVert\right)+\left\lVert(B-D)+(\nu-\beta)I_{m}\right\rVert$
$\leq\left(\sum_{i=1}^{d}\left\lVert A_{i}-C_{i}\right\rVert\right)+\left\lVert B-D\right\rVert+\left(\sum_{i=1}^{d}\left\lVert\lambda_{i}-\alpha_{i}\right\rVert\right)+\left\lVert\nu-\beta\right\rVert$
Lastly we make use of the lipschitz continuity of the smallest singular value of a matrix to get
$\left|\dot{\mu}_{\bm{\lambda}}^{\text{C}}\left(\bm{A},B\right)-\dot{\mu}_{\bm{\alpha}}^{\text{C}}\left(\bm{C},D\right)\right|\leq\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)-L_{(\bm{\alpha},\beta)}\left(\bm{C},D\right)\right\rVert$
@@
$\Lambda_{\epsilon}^{\text{C}}(\bm{A},B),=\bar{\Lambda}_{\epsilon}^{\text{C}}(\bm{A},B)=\dot{\Lambda}_{\epsilon}^{\text{C}}(\bm{A},B).$
The idea is that either the Clifford linear or radial $\epsilon$-pseudospectrum will provide a tuple $(\bm{x},E)$ and a unit eigenstate $\bm{\psi}$ so that (3) holds. As stated in Section 2 we will prove a similar formulation of Lemma 1.2 and 1.3 of [loring]. In other words the $\epsilon$-pseudospectrum gives rise to a unit state that is approximately localized in position and energy.
@@
If $(\bm{\lambda},\nu)$ in $\dot{\Lambda}_{\epsilon}^{\text{C}}\left(\bm{A},B\right)$ for some $\epsilon_{1}\geq 0$ and
@@
then there is a unit state $\bm{\psi}$ such that
@@
$\leq\epsilon^{2}+\left(\sum_{i\not=k}\left\lVert[A_{i},A_{k}]\right\rVert\right)+\lVert F\rVert$
@@
$\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\right)+B^{\dagger}B\right)\bm{\Psi}_{k}\right\rVert\leq\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\otimes I_{2m}\right)+C\right)\bm{\Psi}\right\rVert$
@@
$\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\right)+B^{\dagger}B\right)\bm{\psi}\right\rVert\leq g\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\otimes I_{2m}\right)+C\right)\bm{\Psi}\right\rVert.$
Next we have that
$\sum_{i=1}^{d}\left\lVert A_{i}\bm{\psi}\right\rVert^{2}+\left\lVert B\bm{\psi}\right\rVert^{2}=\left\langle\left(\left(\sum_{i=1}^{d}A_{i}^{2}\right)+B^{\dagger}B\right)\bm{\psi},\bm{\psi}\right\rangle\leq\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\right)+B^{\dagger}B\right)\bm{\psi}\right\rVert.$
This tells completes the proof. ∎
@@
Given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian, $B$ in $\mathcal{M}_{n}(\mathbb{C})$ non-Hermitian. If
$\sqrt{\left(\sum_{i=1}^{d}\left\lVert A_{i}\bm{\psi}-\lambda_{i}\bm{\psi}\right\rVert^{2}\right)+\left\lVert B\bm{\psi}-\nu\bm{\psi}\right\rVert^{2}+\left\lVert B^{\dagger}\bm{\psi}-\overline{\nu}\bm{\psi}\right\rVert^{2}}$ $\leq\epsilon_{1}$
@@
for all $i$, with $F$ as in Equation 10; then $(\bm{\lambda},\nu)$ is in $\Lambda_{\epsilon}(\bm{A},B)$ with $\epsilon=\sqrt{\epsilon_{1}+\epsilon_{2}}$ or $(\bm{\lambda},\nu)$ is in $\Lambda_{\epsilon=0}(\bm{A},B)$.

### v2

@@ 2 Pseudospectrum for joint approximate eigenvectors
Equation (5) is defined to be the quadratic composite operator in Section 1 of [4]. Given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ where each of the $A_{i}$ are Hermitian operators and probe site $\bm{\lambda}$ in $\mathbb{R}^{d+1}$, the quadratic gap is defined to be
@@
The square root is a consequence of proving this quantity is related to the smallest singular value of a related matrix $M_{\bm{\lambda}}(\bm{A})$ which is the topic of Proposition II.1 in [4]. Note that by construction, the quadratic composite operator has real eigenvalues and is positive semidefinite. Informaly a key idea with this gap function is that
$\mu_{\bm{\lambda}}^{\text{C}}(\bm{A})=\sigma_{\min}\left(L_{\bm{\lambda}}\left(\bm{A}\right)\right)\approx\sigma_{\min}\left(\text{Q}_{\bm{\lambda}}\left(\bm{A}\right)\right)\leq\mu_{\bm{\lambda}}^{\text{Q}}\left(\bm{A}\right)$
when $\mu_{\bm{\lambda}}^{Q}(\bm{A})\leq 1$. This also provides a corresponding state $\bm{\psi}$ that is an approximate eigenvector of the constituent observables/matrices. For a comparison between the spectral localizer gap and quadratic gap function, see Proposition II.4 in [4]. The main idea is that the Clifford and quadratic gaps are close if the commutators $A_{i},A_{k}]$ are small, in other words the matrices $A_{i}$ nearly pairwise commute.
3 Non-Hermitian spectral localizer
As mentioned in the introduction, a property of photonic systems is that they exhibit non-Hermitian Hamiltonians. This motivates the question: What happens when $A_{d+1}=B$ is non-Hermitian in the construction of the spectral localizer and in the quadratic composite operator? Or in a physical system setting, what happens when we have a non-Hermitian Hamiltonian $H$ in the construction of the corresponding composite operators? We take inspiration from the constructions and definitions in [3] to motivate this Section. In the process we also introduce a new gap function and explain some of its properties.
One option for a construction of an equivalent spectral localizer is to consider a matrix decomposition of $B$ into the Hermitian part $B_{\textup{H}}$ and its skew-Hermitian part $B_{\textup{S}}$ shown in Equation (6)
$B=B_{\textup{H}}+iB_{\textup{S}},\quad B_{\textup{H}}=(B+B^{\dagger})/2,\quad B_{\textup{S}}=(B-B^{\dagger})/2i.$ (6)
Both $B_{\textup{H}}$ and $B_{\textup{S}}$ are Hermitian, and thus we can use these together with position observables in the non-Hermitian spectral localizer, and thus we would be able to find a unit state $\bm{\psi}$ that is approximately localized in position and energy given the theory above. In particular we could construct
@@ 3 Non-Hermitian spectral localizer
However this goes against the intuition that the non-Hermitian spectral localizer should have similar properties as the Hamiltonian, in this case hermiticity or lack thereof. Also, nothing necessarily stops us from using $B$ in the construction of the spectral localizer in Ref [32, 4] with complex $E$. In other words $L_{(\bm{x},E)}(\bm{X},B)$. This makes it so that the resulting operator is not only not Hermitian, it is also not normal. In fact, we can use the non-normality to show that the (right) eigenvector associated to the smallest eigenvalue ($L_{(\bm{x},E)}(\bm{X},B)\bm{\psi}=\lambda\bm{\psi}$) will give access to an approximate (right) eigenvector of all the matrices $X_{i}$ and $B$. Likewise, a left eigenvector associated with the smallest associated left eigenvalue ($\bm{\phi}^{\dagger}L_{(\bm{x},E)}(\bm{X},B)=\bm{\phi}^{\dagger}\alpha$) will give an approximate left eigenvector of all $X_{i}$ and $H$. However, we will be following the approach in Ref [3] for the construction of the non-Hermitian spectral localizer.
In the construction of the spectral localizer in Section I of [4], an arbitrary Clifford representation can be chosen. However, in the following construction for the non-Hermitian spectral localizer we use a specific Clifford representation. This is because we can choose $d$ even so there will be a Clifford element representation that anticommutes with all the other $d$ elements and is diagonal in a suitable representation. We will use such a representation and replace the diagonal element by a unique choice that allows us to extract information about both left and right eigenvectors. The construction of said Clifford representation is as follows. Note that $\sigma_{x},\sigma_{y},\sigma_{z}$ are the Pauli spin matrices. For $d=2$ we use $\sigma_{x}$ and $\sigma_{y}$ and for $d=1$ we just drop $\sigma_{y}$. For $d=3$ we use $\Gamma_{1}=\sigma_{x},\Gamma_{2}=\sigma_{y},\Gamma_{3}=\sigma_{z}$. The Construction for $d=5$ is $\Gamma_{1}=\sigma_{x}\otimes\sigma_{x}$, $\Gamma_{2}=\sigma_{y}\otimes\sigma_{x}$, $\Gamma_{3}=\sigma_{z}\otimes\sigma_{x}$, $\Gamma_{4}=I_{2}\otimes\sigma_{y}$, $\Gamma_{5}=I_{2}\otimes\sigma_{z}$. Then suppose that $\{\gamma_{i}\}_{i=1}^{d}$ represents a Clifford representation for $d$ odd, we construct a Clifford representation for $d+2$ by computing
@@
If we want the Clifford representation for $d$ even, we just construct the $d+2$ Clifford representation and discard $I_{m}\otimes\sigma_{y}$. We now proceed with the definition of the non-Hermitian spectral localizer as follows.
@@
Given $\bm{A}$ in $\Pi_{i=1}^{d}\mathcal{M}_{n}(\mathbb{C})$ where each of the $A_{i}$ are Hermitian operators, a non-Hermitian operator $B$, and a probe site $(\bm{\lambda},\nu)$ in $\mathbb{R}^{d}\oplus\mathbb{C}$ we define the non-Hermitian spectral localizer to be
@@
Again, this is a special case of the spectral localizer in [3]. From here we can manually compute and determine that the non-Hermitian spectral localizer is not normal. This can be seen below when we compute the product
@@
Despite the non-Hermitian spectral localizer being non-normal, we do have a property that is just as useful if not better. If $B$ is close to being Hermitian then so too is the non-Hermitian spectral localizer. Formally we have that for all $\epsilon>0$, $\left\lVert B-B^{\dagger}\right\rVert<\epsilon$ implies that if we label $L=L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)$, then
@@
$<\epsilon+2\left\lvert\textrm{Im}(\nu)\right\rvert$.
This also tells us that the closeness of the non-Hermitian spectral localizer to Hermitian is also dependent on how big of a value we probe in energy. Further more we have that,
$\left\lVert L^{\dagger}L-LL^{\dagger}\right\rVert$ $\leq\left\lVert(L-L^{\dagger})L\right\rVert+\left\lVert L(L-L^{\dagger})\right\rVert$
$\leq\left\lVert L\right\rVert\left\lVert(L-L^{\dagger})\right\rVert+\left\lVert L\right\rVert\left\lVert(L-L^{\dagger})\right\lVert$
[…]
which tells us that the non-Hermitian spectral localizer being close to hermitian also implies it is close to normal. In a way this is not too surprising from the point of view that if a matrix $A$ is Hermitian, then it follows that $A$ is normal, indeed $A=A^{\dagger}$ implies $A^{\dagger}A=AA=AA^{\dagger}$. To define the gap function for the non-Hermitian spectral localizer, we take the definition in Section 2 of [3] that is used while studying line gapped systems. In particular it is called the spectral localizer gap, but we will refer to it by a different name.
@@
Based on numerical results we conjecture that the Clifford linear gap function is continuous.
Conjecture 3.3.
The Clifford linear gap function is pointwise continuous.
From a topological point of view, especially for line gapped systems, it seems to be instrumental to not look at $\min\left\lvert\textup{Spec}\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)\right\rvert$ but instead to flatten the spectrum and instead look at $\min\left\lvert\textrm{Re}(\textup{Spec}\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)\right\rvert$. When working with non-Hermitian systems, we have to take into consideration both point gapped and line gapped systems. For this reason we propose the following definition for point gapped systems. By numerical computations in Section 7 we will demonstrate that the smallest singular value of the non-Hermitian spectral localizer detects point gaps.
@@
$\dot{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)=\sigma_{\min}\left(L_{\bm{(\lambda},\nu)}\left(\bm{A},B\right)\right).$
A property of the Clifford radial gap is that it’s Lipschitz continuous in a generalized taxi-cab metric. This form of continuity implies the Clifford radial gap is differentiable almost everywhere. This also means that when trying to generate a grid size for numerical simulations, it can be determined ahead of time based on the level of precision desired between data points.
@@
$\left|\dot{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)-\dot{\mu}_{{(\bm{\alpha},\beta)}}^{\text{C}}\left(\bm{C},D\right)\right|\leq 2\left[\left(\sum_{i=1}^{d}\left\lVert A_{i}-C_{i}\right\rVert\right)+\left\lVert B-D\right\rVert+\left(\sum_{i=1}^{d}\left\lvert\lambda_{i}-\alpha_{i}\right\rvert\right)+\left\lvert\nu-\beta\right\rvert\right$].
@@
$\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)-L_{(\bm{\alpha},\beta)}\left(\bm{C},D\right)\right\rVert$ $\leq\sum_{i=1}^{d}\left\lVert\left((A_{i}-C_{i})+(\alpha_{i}-\lambda_{i})I_{2m}\right)\otimes\Gamma_{i}\right\rVert$
$\qquad+\left\lVert\left((B-D)+(\beta-\nu)I_{m}\right)\otimes\begin{bmatrix}I_{m}&\\ &0_{m}\end{bmatrix}\right\rVert$
$\qquad+\left\lVert\left((B-D)^{\dagger}+(\beta-\nu)^{*}I_{m}\right)\otimes\begin{bmatrix}I_{m}&\\ &0_{m}\end{bmatrix}\right\rVert$
$\leq\left(\sum_{i=1}^{d}\left\lVert(A_{i}-C_{i})+(\alpha_{i}-\lambda_{i})I_{2m}\right\rVert\right)+2\left\lVert(B-D)+(\beta-\nu)I_{m}\right\rVert$
$\leq 2\left(\sum_{i=1}^{d}\left\lVert A_{i}-C_{i}\right\rVert\right)+2\left\lVert B-D\right\rVert+2\left(\sum_{i=1}^{d}\left\lvert\lambda_{i}-\alpha_{i}\right\rvert\right)+2\left\lvert\nu-\beta\right\rvert$
Lastly from [27] we know that
$\left\lvert\sigma_{\min}(A)-\sigma_{\min}(B)\right\rvert\leq\left\lVert A-B\right\rVert.$
Thus,
$\left|\dot{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)-\dot{\mu}_{{(\bm{\alpha},\beta)}}^{\text{C}}\left(\bm{C},D\right)\right|\leq\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)-L_{(\bm{\alpha},\beta)}\left(\bm{C},D\right)\right\rVert$
@@
$\Lambda_{\epsilon}^{\text{C}}(\bm{A},B)=\bar{\Lambda}_{\epsilon}^{\text{C}}(\bm{A},B)=\dot{\Lambda}_{\epsilon}^{\text{C}}(\bm{A},B).$
The idea is that either the Clifford linear or radial $\epsilon$-pseudospectrum will provide a tuple $(\bm{\lambda},\nu)$ and a unit eigenvector $\bm{\psi}$ so that (3) holds. As stated in Section 2 we will prove a similar formulation of Lemma 1.2 and 1.3 of [32]. In other words the $\epsilon$-pseudospectrum gives rise to a unit state that is approximately localized in position and energy.
Note that Proposition 3.7 tells us that we can find a state that is an approximate eigenvector of the matrices $A_{i}$ and $B$ or an approximate eigenvector for the matrices $A_{i}$ and $B^{\dagger}$. In other words we either get an approximate right eigenvector for all $A_{i}$, and $B$ or an approximate left eigenvector of all $A_{i}$ and $B$. When $B$ is non-normal, we know the left and right eigenvectors of $B$ are not conjugate to each other, however if B is normal then the left and right eigenvectors will coincide. In any case, it is not possible to find a joint approximate eigenvector for all the matrices $A_{i}$, $B$ and $B^{\dagger}$ without taking a loss in how good the approximate eigenvector is as a left or right eigenvector.
@@
If $(\bm{\lambda},\nu)$ in $\dot{\Lambda}_{\epsilon_{1}}^{\text{C}}\left(\bm{A},B\right)$ for some $\epsilon_{1}\geq 0$ and
@@
for some $\epsilon_{2}\geq 0$ then there is a unit state $\bm{\psi}$ such that either
@@
or
$\sqrt{\left(\sum_{i=1}^{d}\left\lVert A_{i}\bm{\psi}-\lambda_{i}\bm{\psi}\right\rVert^{2}\right)+\left\lVert B^{\dagger}\bm{\psi}-\nu^{*}\bm{\psi}\right\rVert^{2}}\leq\sqrt{2m}\sqrt{\epsilon_{1}^{2}+\epsilon_{2}},$
@@
$\leq\epsilon^{2}+\left(\sum_{i<k}\left\lVert[A_{i},A_{k}]\right\rVert\right)+\lVert F\rVert$
@@
$\left\lVert\left[\left(\sum_{i=1}^{d}A_{i}^{2}\otimes I_{2m}\right)+C\right]\bm{\Psi}\right\rVert\geq\begin{cases}\left\lVert\left[\left(\sum_{i=1}^{d}A_{i}^{2}\right)+B^{\dagger}B\right]\bm{\Psi}_{k}\right\rVert\text{ if }1\leq k\leq m\\ \\ \left\lVert\left[\left(\sum_{i=1}^{d}A_{i}^{2}\right)+BB^{\dagger}\right]\bm{\Psi}_{k}\right\rVert\text{ if }m<k\leq 2m.\end{cases}$
@@
$g\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\otimes I_{2m}\right)+C\right)\bm{\Psi}\right\rVert\geq\begin{cases}\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\right)+B^{\dagger}B\right)\bm{\psi}\right\rVert\text{ if }1\leq k\leq m\\ \\ \left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\right)+BB^{\dagger}\right)\bm{\psi}\right\rVert\text{ if }m<k\leq 2m.\end{cases}$
Next we have that if $1\leq k\leq m$
$\left(\sum_{i=1}^{d}\left\lVert A_{i}\bm{\psi}\right\rVert^{2}\right)+\left\lVert B\bm{\psi}\right\rVert^{2}=\left\langle\left[\left(\sum_{i=1}^{d}A_{i}^{2}\right)+B^{\dagger}B\right]\bm{\psi},\bm{\psi}\right\rangle\leq\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\right)+B^{\dagger}B\right)\bm{\psi}\right\rVert.$
and if $m<k\leq 2m$
$\left(\sum_{i=1}^{d}\left\lVert A_{i}\bm{\psi}\right\rVert^{2}\right)+\left\lVert B^{\dagger}\bm{\psi}\right\rVert^{2}=\left\langle\left[\left(\sum_{i=1}^{d}A_{i}^{2}\right)+BB^{\dagger}\right]\bm{\psi},\bm{\psi}\right\rangle\leq\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\right)+BB^{\dagger}\right)\bm{\psi}\right\rVert.$
This completes the proof. ∎
We will use the following inequality which is a direct consequence of triangle inequality, norms on tensor products as well as Equations (9) and (10); in particular, given a unit vector $\Psi$ we have that
$\quad+\left\lVert\left(\left(\sum_{i<j}\left[A_{i},A_{j}\right]\otimes\Gamma_{i}\Gamma_{j}\right)+F\right)\bm{\Psi}\right\rVert.$ (11)
Lastly note that when the non-Hermitian spectral localizer is invertible,
$\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{-1}\right\rVert^{2}$ $=\left\lVert\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)^{-1}\right\rVert$
$\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{-1}\right\rVert^{-1}$ $=\sqrt{\left\lVert\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)^{-1}\right\rVert^{-1}}$
$=\sqrt{\sigma_{\min}\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)}$
This implies that
$\dot{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)$ $=\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{-1}\right\rVert^{-1}$
$=\sqrt{\min_{\left\lVert\mathbf{v}\right\rVert=1}\left\lVert\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)\mathbf{v}\right\rVert}$
$\leq\sqrt{\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\bm{\Psi}\right\rVert}.$ (12)
Proposition 3.8 states that if we have a unit state that is approximately a right left eigenvector of $A_{i}$ and $B$ then the corresponding tuple of eigenvalues lie in some $\epsilon$-pseudospectrum.
@@
Suppose that $\bm{A}$ is in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian, and further suppose $B$ is in $\mathcal{M}_{n}(\mathbb{C})$ and is non-Hermitian. If there is a unit state $\bm{\psi}$ such that
$2\left(\sum_{i=1}^{d}\left\lVert A_{i}\bm{\psi}-\lambda_{i}\bm{\psi}\right\rVert^{2}\right)$ $+\left\lVert B\bm{\psi}-\nu\bm{\psi}\right\rVert^{2}+\left\lVert B^{\dagger}\bm{\psi}-\nu^{*}\bm{\psi}\right\rVert^{2}\leq\epsilon_{1}$
and
@@
for all $i$, with $F$ as in Equation 10; then $(\bm{\lambda},\nu)$ is in $\dot{\Lambda}_{\epsilon}(\bm{A},B)$ with $\epsilon=\sqrt{\frac{1}{\sqrt{2}}\epsilon_{1}+\epsilon_{2}}$ or $(\bm{\lambda},\nu)$ is in $\dot{\Lambda}_{\epsilon=0}(\bm{A},B)$.